·て 2. (13 points) Consider the following demand and supply functions: Qd=20-2p and Q-4t2P a. Find the equilibrium price and quantity. Show your work. (2) b. Find consumer and producer surplus at the equilibrium. Show your work. (4) c. If a tax of $1 is imposed on the seller, what will be the new equilibrium prices (for buyer an seller) and the quantity? Show your work. (5) Find the DWL created by the tax. Show your work. (2) d.
Answer #1

a). Solution :- At an equilibrium, Demand = Supply.

20 - 2P = (-) 4 + 2P

20 + 4 = 2P + 2P

24 = 4P

P = 24 / 4

P = $ 6.

Q = 20 - 2 * 6 (Put the value of P = 6 in the demand equation).

Q = 20 - 12

Q = 8 units.

Conclusion :-

Equilibrium price (P) $ 6.
Equilibrium quantity (Q) 8 units.

b). Solution :- Consumer surplus = 0.50 * Base * Height.

= 0.50 * (10 - 6) * (8 - 0)

= 0.50 * 4 * 8

= $ 16.

Producer surplus = 0.50 * Base * Height.

= 0.50 * (6 - 2) * (8 - 0)

= 0.50 * 4 * 8

= $ 16.

Conclusion :-

Consumer surplus $ 16.
Producer surplus $ 16.